OpenStudy (anonymous):

Please help! Find the following infinite sum, if its exists. If one does not exist enter "no infinite sum." (Look at response.)

OpenStudy (anonymous):

\[\sum_{n-1}^{\infty}4(1/3)^(n-1)\]

OpenStudy (anonymous):

n-1 is raised (^)

OpenStudy (anonymous):

\[4\sum _{n=1}^{\infty } \left(\frac{1}{3}\right)^{n-1}\]

OpenStudy (anonymous):

I am going to change it

OpenStudy (anonymous):

\[4\sum _{n=0}^{\infty } \left(\frac{1}{3}\right)^n\]

OpenStudy (anonymous):

That's correct :/ Sorry

OpenStudy (anonymous):

This is formula \[\sum _{n=0}^{\infty } a^n =\text{ }\frac{1}{1-r}\] \[\sum _{n=0}^{\infty } (1/3)^n =\text{ }\frac{1}{1-(1/3)}\]

OpenStudy (anonymous):

1/(2/3) 3/2

OpenStudy (anonymous):

So the infinite sum would be 3/2?

OpenStudy (anonymous):

yes, by the way we were able to do it because 1/3 is less than 1

OpenStudy (zarkon):

times 4

OpenStudy (anonymous):

Huh Zarkon?

OpenStudy (zarkon):

3/2*4=6

OpenStudy (anonymous):

oh yeah , we forgot about 4 infront

OpenStudy (anonymous):

Oh I see. So final answer= 6. Thank you guys! :)